Conversions 1 in. = 2.54 cm.
a) How many square centimeters are in a square inch?
b) How many cubic centimeters are in a cubic inch?
c) How long is a centimeter in terms of inches?

Answer by mananth(16949)

Conversions 1 in. = 2.54 cm.
a) How many square centimeters are in a square inch?
1 sq.in = 1"*1" = 2.54*2.54=6.451 cm^2
b) How many cubic centimeters are in a cubic inch?
1 in^3 = 1*1*1 = 2.54*2.54*2.54=16.387 cm^3
c) How long is a centimeter in terms of inches?
1/2.54= 0.393 in
...
m.ananth@hotmail.ca